# Gravitational wave probes of particle dark matter: a review

### v4

@@ III.2.2 KAGRA
KAGRA offers unique opportunities to search for DM. In particular, the different materials used in KAGRA’s mirrors result in significantly different charge-to-mass ratios for baryon-lepton number, enhancing the sensitivity to vector DM coupling to this quantum number Michimura et al. (2020). While LIGO and Virgo currently set the strongest limits on dark photon coupling to baryon-lepton number – obtainable from the results in Fig. 4 by multiplying the limits on baryon number coupling by four Abbott et al. (2022c) – this difference in mirror materials means that once KAGRA reaches comparable sensitivity levels, it will outperform LIGO in constraining such couplings Michimura et al. (2020).
@@
The raw outputs of PRCL and MICH correspond to time-varying displacements of the form Abac et al. (2024):
@@
where $l_{x}=26.7\,\text{m}$ and $l_{y}=23.3\,\text{m}$ are the distances between the beam splitter and the input mirrors, and $l_{p}=41.6\,\text{m}$ is the length of the power recycling cavity. All these lengths are much shorter than the DM coherence length $L_{\text{coh}}$. See Fig. 1 of Abac et al. (2024) for a schematic of the interferometer.
No significant signals were found in this analysis of KAGRA O3 data Abac et al. (2024), and upper limits on the coupling strength were derived following the procedure outlined in Section III.1.4. These limits, shown in Fig. 6 for each channel, exhibit many narrow-band noise artifacts that limit sensitivity at low frequencies. Although not yet competitive with LIGO constraints, future KAGRA searches are expected to provide powerful limits on this type of DM due to the pronounced difference in baryon-lepton charge-to-mass ratios among its mirrors Michimura et al. (2020).
Figure 6: Taken from Abac et al. (2024). 95% confidence-level upper limit on the strength of the coupling between dark photons and baryon-lepton number using KAGRA O3 data and the method described in Section III.1.4, derived from two particular interferometer channels: MICH (blue line) and PRCL (orange line). The physics behind this form of DM has been discussed in Section II.4. Because the KAGRA input mirrors, end mirrors, and power recycling mirrors are made of different material, the coupling strength of dark photons to baryon-lepton number is enhanced (Eq. 54), and appears more strongly in these particular channels than in the ordinary strain one. In the low-mass (low-frequency) range, many narrow-band noise disturbances of unknown origin are seen.
@@ III.2.3 LISA Pathfinder
Such a search was preformed using LISA Pathfinder data Miller and Mendes (2023) using the excess power method described in Section III.1.2 for ultralight dark photons coupling to baryons, however, no physical constraints on the coupling strength were obtained, since the noise level was too high, the arm length was too short (only $40$ cm compared to the expected LISA arm lengths of $\mathcal{O}(10^{6})$ km), and only one arm existed. However, the search procedure designed will serve as a roadmap for future analyses of spaced-based detector data, since many problems, including limited sampling at low frequency, gaps, noise non-stationarities and glitches had to be handled in this analysis of data from a space-based detector.
@@
After this search was performed, however, it was realized that data from a different channel would produce tighter constraints, particularly in the case of baryon-lepton coupling. Here, the relative acceleration of the spacecraft and one of the test masses would produce a much stronger signal than that which was searched for previously because the charge-to-mass ratios of each are different Frerick et al. (2024). This echoes the logic presented in Section III.2.2 to use KAGRA to search for dark photons coupling to baryon-lepton number.
@@
A conservative upper limit on the baryon-lepton coupling was set without performing a search, assuming no signal would have been detected in LISA Pathfinder data, as shown in Fig. 7, by using a reference acceleration power spectral density of the spacecraft with respect to one of the test masses. We can see that the upper limits from the relative acceleration of space-craft/ test mass are much better than that from Miller and Mendes (2023), labeled (decoherence). The constraint on B-L coupling is stronger with respect to that of baryon coupling because the test mass and spacecraft are primarily made of gold, and gold/carbon, respectively.
@@
Figure 7: Taken from Frerick et al. (2024). Upper limits on the coupling of dark photons to baryon-lepton number considering the relative acceleration of the spacecraft and one of the test masses in LISA Pathfinder (dark blue region). The physics behind this form of DM has been discussed in Section II.4, and constrains the same model as in Figs. 6 and 4, though at much lower boson masses. The constraints derived from a search for the induced acceleration of the two test masses by dark photons in LISA Pathfinder data is shown as “decoherence” (dark red region). Using the relative acceleration caused by dark photon interaction with baryon-lepton number allows for stringent constraints to be placed on the coupling, since the spacecraft and test masses are made of different materials. Constraints from other experiments (Eöt-Wash and MICROSCOPE) and future projections for LISA, MAGIS-100 and torision balance are shown as well. Details on the curves coming from STE-QUEST and Asteroids can be found in O’HARE (2020).
@@
Only B-L constraints are shown from Frerick et al. (2024) because those arising from coupling to baryons are not yet strong enough to probe a tighter constraint than that provided by MICROSCOPE. However, many conservative assumptions were made in Frerick et al. (2024), including the composition of the spacecraft and the use of a reference acceleration power spectral density, both of which could, if studied in more detail, lead to more stringent constraints for both the baryon and B-L coupling scenarios.
@@ III.3 Pulsar timing arrays
As described in Section II.3, ultralight scalar DM can also induce variations in pulse ToAs Khmelnitsky and Rubakov (2014), though the relevant mass range lies many orders of magnitude below that probed by ground-based interferometers. Such scalar fields can cause coherent oscillations in the gravitational potential or effective constants of nature, and may also generate stochastic perturbations in the metric Kim and Mitridate (2024), which could lead to further timing variations over long baselines Kim (2023).
@@ III.3.1 Gravitationally-interacting ultralight dark matter
Figure 8: Taken from Smarra et al. (2023). 95% confidence-level upper limits from analyses of European pulsar timing array data on ultralight dark matter interacting only gravitationally and changing the ToAs of pulses in an analogous way that GWs do. The physics behind this form of DM has been discussed in Section II.6. The panels show (a) the dimensionless strain amplitude $\Psi_{\text{c}}$ and (b) the fraction of the DM energy density that scalar DM could compose as a function of DM mass and frequency. $\rho_{\text{DM}}=0.4\text{GeV}/\text{cm}^{3}$. For comparison, upper limits on these quantities are shown from previous searches (PPTA Bayes 2018, NANOGrav 2018) Porayko and Postnov (2014); Porayko et al. (2018). Furthermore, the dashed line indicates the signal amplitude assuming that this form of DM comprises all of DM; thus, curves that lie above this dashed line do not represent physical constraints. Upper limits on the amplitude are shown for the three cases outlined in Section III.3.1. In the right panel, we zoom in on the excluded ultralight dark matter masses. The horizontal dotted line represents the scenario in which ultralight dark matter is responsible for all of the local DM energy density. The right panel shows that scalar ultralight dark-matter particles with mass $24.0<\text{log}_{10}\text{~}(m_{\rm DM}/\text{eV})<-23.7$ can comprise at most $30-40\%$ of $\rho_{\text{DM}}$, while particles with masses $23.7<\text{log}_{10}\text{~}(m_{\rm DM}/\text{eV})<-23.3$ could comprise up to $\sim 70$ % of $\rho_{\text{DM}}$.
Additionally, NANOGrav data has been be used to constrain this kind of ultralight dark matter Afzal et al. (2023), though the constraints are weaker than from the EPTA. In particular, NANOGrav cannot yet probe a physical constraint for the fraction of DM energy density that this model of DM could compose, likely because their dataset only spans 15 years, while the European pulsar timing array collaboration analyzed almost 25 years of data.
@@ III.3.2 Constraints on scalar conformal dark matter
While the previous subsection considered generic strain-like gravitational interactions of scalar ultralight dark matter, scalar conformal couplings represent a distinct class of models in which ULDM interacts through the trace of the energy–momentum tensor, As discussed in Section II.6. This model of DM can also be constrained with pulsar timing array data, and in particular Ref. Smarra et al. (2024) uses European pulsar timing array data to constrain both linear (FJBD-type) and quadratic (DEF-type) conformal couplings.
In Fig. 9, we show the upper limits on the linear coupling parameter $\alpha$ using European pulsar timing array data. However, despite showing upper limits (that assume no signal), Ref. Smarra et al. (2024) notes that there is some evidence for additional signal power at two masses: $m\sim 10^{-22.7}$ eV and $m\sim 10^{-21.4}$ eV, though other physical processes could be responsible for these excesses. It is worth noting that the effective change in the neutron-star moment of inertia in this scenario arises from the universal rescaling of the metric due to the conformal coupling, and is therefore distinct from the composition-dependent moment of inertia variations discussed in Section II.3, where scalar fields couple directly to quarks and nucleons. The constraints arising from the interaction of dilatons with standard-model particles in neutron stars will be discussed in the next subsection.
Figure 9: Taken from Smarra et al. (2024). Constraints on the coupling $\log_{10}\alpha$ of conformal DM to gravity at the 95% confidence-level (solid line) for the correlated case compared to the analytical estimate obtained by using the upper limits from the European pulsar timing array search in Smarra et al. (2024) to estimate $\alpha$ (brown dashed lines). The physics behind this form of DM has been discussed in Section II.6. $\rho_{\text{DM}}=0.4\,\text{GeV/cm}^{3}$. Here, this model of DM would change the moment of inertia of neutron stars used in pulsar timing arrays, resulting in changes of the ToA of the pulses. If conformal DM does not comprise all of DM, the limits will weaken, which is shown by the dotted line, for values of $\rho$ given in Smarra et al. (2023). AP4 indicates that these limits are derived assuming the AP4 equation-of-state of neutron stars Akmal et al. (1998).
@@ III.3.3 Variation of fundamental constants constraints
The strongest constraints from pulsar timing arrays are on the couplings to the electron and muons. This is in contrast to constraints that come from atomic clock experiments, which also aim to find ultralight dark matter but are insensitive to the electron coupling, since changing the electron mass does not affect the spacing between energy levels Kaplan et al. (2022). In terms of (the lack of) other constraints on muons, laboratory experiments do not have enough muons to harness on earth to study the coupling of DM to them, while, in contrast, neutron stars have an abundance of muons.
In Fig. 10, there is a gray-shaded region that indicates the situation in which $A<A_{\rm grav}$, where $A$ is given by the amplitude of Eq. 40 or Eq. 44. In other words, these constraints on ultralight dark matter coupling to standard-model particles are valid when the coupling of DM to gravity can be neglected, i.e. when Kaplan et al. (2022):
@@
Figure 10: Taken from Afzal et al. (2023). Upper limits from an analysis of NANOGrav data on the coupling of ultralight dark matter to various standard-model particles: quarks, electrons, muons, photons and gluons, shown in red for the correlated case, and red dashed for the uncorrelated case. In this model, DM would alter the moment of inertia of neutron stars and thus affect the ToAs of pulses in pulsar timing arrays. The physics behind this form of DM has been discussed in Section II.3.2. The constraints on one coupling constant assume that all other coupling constants are set to 0. The black dashed line, and the gray shaded region, indicate where the amplitude of the signal would be less than the amplitude of the signal arising from DM coupling only to gravity. Current constraints “Rb/Cs atomic clocks” (purple) are from Hees et al. (2016), “Al/Hg atomic clocks” (turquoise) are from Beloy et al. (2021), “MICROSCOPE” (teal) are from Bergé et al. (2018), “H/Si clock shift” (orange) are from Beloy et al. (2021), and “NS binary system” are from Kumar Poddar et al. (2019) and Dror et al. (2020). The shape of these constraints is similar to those in Fig. 11, since both searches look for nearly sinusoidal signals embedded in pulsar timing array data with similar noise characteristics, even though the underlying DM coupling mechanisms differ.
@@ III.3.4 Constraints on vector dark matter
Figure 11: Taken from Afzal et al. (2023). Upper limits from analyses of NANOGrav pulsar timing array data on the coupling of DM to baryon and baryon-lepton number, shown for the correlated (solid lines) and uncorrelated (dashed lines) cases. In this scenario, dark photon DM couples to standard model particles in both the Earth and the pulsars. The physics behind this form of DM has been discussed in Section II.4. The black dashed line and gray shaded region indicate where the amplitude of the signal would be less than that from gravitational coupling alone. Constraints from equivalence principle tests (“MICROSCOPE”) are shown in teal (Bergé et al., 2018), and previous limits from the PPTA Collaboration are shown in yellow (Xue et al., 2022). The similarity in shape between the curves in this figure and those in Fig. 10 arises because both searches target nearly sinusoidal signals embedded in pulsar timing array data with similar noise characteristics. Although the underlying DM models differ—scalar versus vector, moment of inertia changes versus direct forces—the observable effect in each case is the same: modulations in pulse ToAs. The coupling strength to standard-model particles determines the amplitude of these signals.
Furthermore, recent evidence for a stochastic GW background reported by NANOGrav has also been interpreted in terms of vector ultralight dark matter. In particular, Ref. Chowdhury et al. (2024) showed that a vector ultralight dark matter model coupled to muons ($L_{\mu}-L_{\tau}$) can provide an even better fit to NANOGrav’s observations than a conventional stochastic GW background. In this scenario, pulsars undergo oscillations due to their muon content, leading to a spin-dependent frequency shift in the pulses. This produces angular correlations that differ from the standard Hellings–Downs curve expected for a stochastic GW background, thereby offering a way to test the vector ultralight dark-matter hypothesis against future pulsar timing array data.

### v5

@@ III.2.2 KAGRA
KAGRA offers unique opportunities to search for DM. In particular, the different materials used in KAGRA’s mirrors result in significantly different charge-to-mass ratios for baryon-lepton number, enhancing the sensitivity to vector DM coupling to this quantum number Michimura:2020vxn. While LIGO and Virgo currently set the strongest limits on dark photon coupling to baryon-lepton number – obtainable from the results in Fig. 4 by multiplying the limits on baryon number coupling by four LIGOScientific:2021odm – this difference in mirror materials means that once KAGRA reaches comparable sensitivity levels, it will outperform LIGO in constraining such couplings Michimura:2020vxn.
@@
The raw outputs of PRCL and MICH correspond to time-varying displacements of the form KAGRA:2024ipf:
@@
where $l_{x}=26.7\,\text{m}$ and $l_{y}=23.3\,\text{m}$ are the distances between the beam splitter and the input mirrors, and $l_{p}=41.6\,\text{m}$ is the length of the power recycling cavity. All these lengths are much shorter than the DM coherence length $L_{\text{coh}}$. See Fig. 1 of KAGRA:2024ipf for a schematic of the interferometer.
No significant signals were found in this analysis of KAGRA O3 data KAGRA:2024ipf, and upper limits on the coupling strength were derived following the procedure outlined in Section III.1.4. These limits, shown in Fig. 6 for each channel, exhibit many narrow-band noise artifacts that limit sensitivity at low frequencies. Although not yet competitive with LIGO constraints, future KAGRA searches are expected to provide powerful limits on this type of DM due to the pronounced difference in baryon-lepton charge-to-mass ratios among its mirrors Michimura:2020vxn.
Figure 6: Taken from KAGRA:2024ipf. 95% confidence-level upper limit on the strength of the coupling between dark photons and baryon-lepton number using KAGRA O3 data and the method described in Section III.1.4, derived from two particular interferometer channels: MICH (blue line) and PRCL (orange line). The physics behind this form of DM has been discussed in Section II.4. Because the KAGRA input mirrors, end mirrors, and power recycling mirrors are made of different material, the coupling strength of dark photons to baryon-lepton number is enhanced (Eq. 54), and appears more strongly in these particular channels than in the ordinary strain one. In the low-mass (low-frequency) range, many narrow-band noise disturbances of unknown origin are seen.
@@ III.2.3 LISA Pathfinder
Such a search was preformed using LISA Pathfinder data Miller:2023kkd using the BSD excess power method described in Section III.1.2 for ultralight dark photons coupling to baryons, however, no physical constraints on the coupling strength were obtained, since the noise level was too high, the arm length was too short (only $40$ cm compared to the expected LISA arm lengths of $\mathcal{O}(10^{6})$ km), and only one arm existed. However, the search procedure designed will serve as a roadmap for future analyses of spaced-based detector data, since many problems, including limited sampling at low frequency, gaps, noise non-stationarities and glitches had to be handled in this analysis of data from a space-based detector.
@@
After this search was performed, however, it was realized that data from a different channel would produce tighter constraints, particularly in the case of baryon-lepton coupling. Here, the relative acceleration of the spacecraft and one of the test masses would produce a much stronger signal than that which was searched for previously because the charge-to-mass ratios of each are different Frerick:2023xnf. This echoes the logic presented in Section III.2.2 to use KAGRA to search for dark photons coupling to baryon-lepton number.
@@
A conservative upper limit on the baryon-lepton coupling was set without performing a search, assuming no signal would have been detected in LISA Pathfinder data, as shown in Fig. 7, by using a reference acceleration power spectral density of the spacecraft with respect to one of the test masses. We can see that the upper limits from the relative acceleration of space-craft/ test mass are much better than that from Miller:2023kkd, labeled (decoherence). The constraint on B-L coupling is stronger with respect to that of baryon coupling because the test mass and spacecraft are primarily made of gold, and gold/carbon, respectively.
@@
Figure 7: Taken from Frerick:2023xnf. Upper limits on the coupling of dark photons to baryon-lepton number considering the relative acceleration of the spacecraft and one of the test masses in LISA Pathfinder (dark blue region). The physics behind this form of DM has been discussed in Section II.4, and constrains the same model as in Figs. 6 and 4, though at much lower boson masses. The constraints derived from a search for the induced acceleration of the two test masses by dark photons in LISA Pathfinder data is shown as “decoherence” (dark red region). Using the relative acceleration caused by dark photon interaction with baryon-lepton number allows for stringent constraints to be placed on the coupling, since the spacecraft and test masses are made of different materials. Constraints from other experiments (Eöt-Wash and MICROSCOPE) and future projections for LISA, MAGIS-100 and torision balance are shown as well. Details on the curves coming from STE-QUEST and Asteroids can be found in ciaran_o_hare_2020_3932430.
@@
Only B-L constraints are shown from Frerick:2023xnf because those arising from coupling to baryons are not yet strong enough to probe a tighter constraint than that provided by MICROSCOPE. However, many conservative assumptions were made in Frerick:2023xnf, including the composition of the spacecraft and the use of a reference acceleration power spectral density, both of which could, if studied in more detail, lead to more stringent constraints for both the baryon and B-L coupling scenarios.
@@ III.3 Pulsar timing arrays
As described in Section II.3, ultralight scalar DM can also induce variations in pulse ToAs Khmelnitsky:2013lxt, though the relevant mass range lies many orders of magnitude below that probed by ground-based interferometers. Such scalar fields can cause coherent oscillations in the gravitational potential or effective constants of nature, and may also generate stochastic perturbations in the metric Kim:2023kyy, which could lead to further timing variations over long baselines Kim:2023pkx.
@@ III.3.1 Gravitationally-interacting ultralight dark matter
Figure 8: Taken from EuropeanPulsarTimingArray:2023egv. 95% confidence-level upper limits from analyses of European pulsar timing array data on ultralight dark matter interacting only gravitationally and changing the ToAs of pulses in an analogous way that GWs do. The physics behind this form of DM has been discussed in Section II.6. The panels show (a) the dimensionless strain amplitude $\Psi_{\text{c}}$ and (b) the fraction of the DM energy density that scalar DM could compose as a function of DM mass and frequency. $\rho_{\text{DM}}=0.4\text{GeV}/\text{cm}^{3}$. For comparison, upper limits on these quantities are shown from previous searches (PPTA Bayes 2018, NANOGrav 2018) Porayko:2014rfa; Porayko:2018sfa. Furthermore, the dashed line indicates the signal amplitude assuming that this form of DM comprises all of DM; thus, curves that lie above this dashed line do not represent physical constraints. Upper limits on the amplitude are shown for the three cases outlined in Section III.3.1. In the right panel, we zoom in on the excluded ultralight dark matter masses. The horizontal dotted line represents the scenario in which ultralight dark matter is responsible for all of the local DM energy density. The right panel shows that scalar ultralight dark-matter particles with mass $24.0<\text{log}_{10}\text{~}(m_{\rm DM}/\text{eV})<-23.7$ can comprise at most $30-40\%$ of $\rho_{\text{DM}}$, while particles with masses $23.7<\text{log}_{10}\text{~}(m_{\rm DM}/\text{eV})<-23.3$ could comprise up to $\sim 70$ % of $\rho_{\text{DM}}$.
Additionally, NANOGrav data has been be used to constrain this kind of ultralight dark matter NANOGrav:2023hvm, though the constraints are weaker than from the EPTA. In particular, NANOGrav cannot yet probe a physical constraint for the fraction of DM energy density that this model of DM could compose, likely because their dataset only spans 15 years, while the European pulsar timing array collaboration analyzed almost 25 years of data.
@@ III.3.2 Constraints on scalar conformal dark matter
While the previous subsection considered generic strain-like gravitational interactions of scalar ultralight dark matter, scalar conformal couplings represent a distinct class of models in which ULDM interacts through the trace of the energy–momentum tensor, As discussed in Section II.6. This model of DM can also be constrained with pulsar timing array data, and in particular Ref. Smarra:2024kvv uses European pulsar timing array data to constrain both linear (FJBD-type) and quadratic (DEF-type) conformal couplings.
In Fig. 9, we show the upper limits on the linear coupling parameter $\alpha$ using European pulsar timing array data. However, despite showing upper limits (that assume no signal), Ref. Smarra:2024kvv notes that there is some evidence for additional signal power at two masses: $m\sim 10^{-22.7}$ eV and $m\sim 10^{-21.4}$ eV, though other physical processes could be responsible for these excesses. It is worth noting that the effective change in the neutron-star moment of inertia in this scenario arises from the universal rescaling of the metric due to the conformal coupling, and is therefore distinct from the composition-dependent moment of inertia variations discussed in Section II.3, where scalar fields couple directly to quarks and nucleons. The constraints arising from the interaction of dilatons with standard-model particles in neutron stars will be discussed in the next subsection.
Figure 9: Taken from Smarra:2024kvv. Constraints on the coupling $\log_{10}\alpha$ of conformal DM to gravity at the 95% confidence-level (solid line) for the correlated case compared to the analytical estimate obtained by using the upper limits from the European pulsar timing array search in Smarra:2024kvv to estimate $\alpha$ (brown dashed lines). The physics behind this form of DM has been discussed in Section II.6. $\rho_{\text{DM}}=0.4\,\text{GeV/cm}^{3}$. Here, this model of DM would change the moment of inertia of neutron stars used in pulsar timing arrays, resulting in changes of the ToA of the pulses. If conformal DM does not comprise all of DM, the limits will weaken, which is shown by the dotted line, for values of $\rho$ given in EuropeanPulsarTimingArray:2023egv. AP4 indicates that these limits are derived assuming the AP4 equation-of-state of neutron stars Akmal:1998cf.
@@ III.3.3 Variation of fundamental constants constraints
The strongest constraints from pulsar timing arrays are on the couplings to the electron and muons. This is in contrast to constraints that come from atomic clock experiments, which also aim to find ultralight dark matter but are insensitive to the electron coupling, since changing the electron mass does not affect the spacing between energy levels Kaplan:2022lmz. In terms of (the lack of) other constraints on muons, laboratory experiments do not have enough muons to harness on earth to study the coupling of DM to them, while, in contrast, neutron stars have an abundance of muons.
In Fig. 10, there is a gray-shaded region that indicates the situation in which $A<A_{\rm grav}$, where $A$ is given by the amplitude of Eq. 40 or Eq. 44. In other words, these constraints on ultralight dark matter coupling to standard-model particles are valid when the coupling of DM to gravity can be neglected, i.e. when Kaplan:2022lmz:
@@
Figure 10: Taken from NANOGrav:2023hvm. Upper limits from an analysis of NANOGrav data on the coupling of ultralight dark matter to various standard-model particles: quarks, electrons, muons, photons and gluons, shown in red for the correlated case, and red dashed for the uncorrelated case. In this model, DM would alter the moment of inertia of neutron stars and thus affect the ToAs of pulses in pulsar timing arrays. The physics behind this form of DM has been discussed in Section II.3.2. The constraints on one coupling constant assume that all other coupling constants are set to 0. The black dashed line, and the gray shaded region, indicate where the amplitude of the signal would be less than the amplitude of the signal arising from DM coupling only to gravity. Current constraints “Rb/Cs atomic clocks” (purple) are from Hees:2016gop, “Al/Hg atomic clocks” (turquoise) are from BACON:2020ubh, “MICROSCOPE” (teal) are from Berge:2017ovy, “H/Si clock shift” (orange) are from BACON:2020ubh, and “NS binary system” are from KumarPoddar:2019ceq and Dror:2019uea. The shape of these constraints is similar to those in Fig. 11, since both searches look for nearly sinusoidal signals embedded in pulsar timing array data with similar noise characteristics, even though the underlying DM coupling mechanisms differ.
@@ III.3.4 Constraints on vector dark matter
Figure 11: Taken from NANOGrav:2023hvm. Upper limits from analyses of NANOGrav pulsar timing array data on the coupling of DM to baryon and baryon-lepton number, shown for the correlated (solid lines) and uncorrelated (dashed lines) cases. In this scenario, dark photon DM couples to standard model particles in both the Earth and the pulsars. The physics behind this form of DM has been discussed in Section II.4. The black dashed line and gray shaded region indicate where the amplitude of the signal would be less than that from gravitational coupling alone. Constraints from equivalence principle tests (“MICROSCOPE”) are shown in teal (Berge:2017ovy), and previous limits from the PPTA Collaboration are shown in yellow (PPTA:2021uzb). The similarity in shape between the curves in this figure and those in Fig. 10 arises because both searches target nearly sinusoidal signals embedded in pulsar timing array data with similar noise characteristics. Although the underlying DM models differ—scalar versus vector, moment of inertia changes versus direct forces—the observable effect in each case is the same: modulations in pulse ToAs. The coupling strength to standard-model particles determines the amplitude of these signals.
Furthermore, recent evidence for a stochastic GW background reported by NANOGrav has also been interpreted in terms of vector ultralight dark matter. In particular, Ref. Chowdhury:2023xvy showed that a vector ultralight dark matter model coupled to muons ($L_{\mu}-L_{\tau}$) can provide an even better fit to NANOGrav’s observations than a conventional stochastic GW background. In this scenario, pulsars undergo oscillations due to their muon content, leading to a spin-dependent frequency shift in the pulses. This produces angular correlations that differ from the standard Hellings–Downs curve expected for a stochastic GW background, thereby offering a way to test the vector ultralight dark-matter hypothesis against future pulsar timing array data.